OpenStudy (anonymous):

Sixty-three is 90% of what number?

OpenStudy (anonymous):

@Saylilbaby

OpenStudy (anonymous):

you just have to solve this for x: \[\frac{90}{100}x=63\]

OpenStudy (anonymous):

70

OpenStudy (anonymous):

which one is right i need to know

OpenStudy (anonymous):

its 70

OpenStudy (anonymous):

70 is the answer to the equation i posted, so both are right!

OpenStudy (mathstudent55):

Sixty-three is 90% of what number? 63 = 90/100 * x \(\dfrac{90x}{100} = 63\) \(90x = 6300\) x = 70 Answer: 70%

OpenStudy (anonymous):

i was trying to show you the equation that represents the situation stated in the problem....

OpenStudy (anonymous):

@mathstudent55 actually the answer is the number 70, not 70%

OpenStudy (mathstudent55):

Correct. The answer is 70, not 70%. Thanks for pointing it out.